• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:02:38 on November13,2002

    Entry number 90041

    keyword=Shift Summary

    Day shift 11/13/02
    John LeRose (SL)
    Ron Gilman (TO)
    Rakhsha Nasseripour

    Production running Q3_j40
    Left arm: 3.783 GeV/c @ 24.830 degrees
    Right arm: 1.8535 GeV/c @ 48.660 degrees

    At start of shift run 3141 in progress (since 07:49)
    08:50 start 3142
    09:40 start 3143
    10:06 ended 3143 Thought beam was going away for beam studies but then there was a 15-minute delay
    10:08 started 3144
    10:25 ended 3144 Beginning of beam Studies (short controlled access for Bogdan)
    11:15 Bogdan is out, back to beam permit. MCC requests we remove the target for BCM calibrations
    BCM calibrations done runs 3146 and 3147 BcmLog_72 and BcmLog_73 (MCC has some trouble getting stable current at the very low current settings)
    14:03 back in production mode run 3148 started
    15:03 run 3149 started still running at shift end